K2SO4 is a strong electrolyte. Determine the concentration of each of the individual ions in a 0.700 M K2SO4 solution.
Answer #1

K2SO4 is a strong electrolyte. gives 100% dissociation

so that ,

Concentration of K^+   = 0.7*2 = 1.4 M

Concentration of So4^2- = 0.7 M